Using a long rod that has length μ, you are going to lay out asquare plot in which the length of each side is μ. Thus thearea of the plot will be μ2. However, you do notknow the value of μ, so you decide to make n independentmeasurements X1, X2, ..., Xn ofthe length. Assume that each Xi has mean μ(unbiased measurements) and varianceσ2.
a. Show that x 2 is not an unbiased estimator for μ2.
b. For what value k is the estimator x 2 -kS2 unbiased for μ2?
Answer #1

We know that \(E(\bar{X})=\mu\) \(V(X)=\frac{\sigma^{2}}{n}\)

and \(V(X)=E\left(X^{2}\right)-[E(X)]^{2}\)

\(\Rightarrow E\left(X^{2}\right)=V(X)+[E(X)]^{2}\)

\(\therefore E\left(\bar{X}^{2}\right)=V(\bar{X})+[E(\bar{X})]^{2}\)

\(\quad=\frac{\sigma^{2}}{n}+\mu^{2}\)

\(\neq \mu\)

THerefore \(\bar{X}^{2}\) is not an unbiased estimator for \(\mu^{2}\)

(b) Now, E \(\left[\bar{X}^{2}-K S^{2}\right]=\mathrm{E}\left(\bar{X}^{2}\right)-k E\left(S^{2}\right)\)

\(=\frac{\sigma^{2}}{n}+\mu^{2}-k \sigma^{2}\)

For, \(\bar{X}^{2}-K S^{2}\) to be an unbiased estimate for \(\mu^{2}\) \(E\left[\bar{X}^{2}-K S^{2}\right]=\mu^{2}\)

Thus, \(\frac{\sigma^{2}}{n}+\mu^{2}-k \sigma^{2}=\mu^{2}\)

\(\Rightarrow \frac{\sigma^{2}}{n}=k \sigma^{2}\)

\(\Rightarrow k=\frac{1}{n}\)